two cars leave an intersection. One car travela north the other east. When the car traveling north had gone 9 mi, the distance between the cars was 3 mi more than the distance traveled by the car heading east. How far had the eastbond car travled?
Answer by mananth(16949)

let east bound car travel x miles
North bound - 9 miles
At right angles
Use pythagoras theorem
x^2++9^2=(x+3)^2
x^2+81=x^2+6x+9
6x=72
x=12